What is the Nernst (equilibrium) potential of a monovalent cation (+ valence) at 37oC given the intracellular ion concentration is 10 mM and the extracellular ion concentration is 80 mM?
Answer #1

According to Nernst equation, nernst potential can be derived by,

Veq = (RT/zF)* ln(Xout/Xin)

Where,

Veq - Nernst potential

R - Universal gasconstant - 8.314 J/K/mol

T- Temperature in Kelvin (K= 0C+ 373.15) = Here 37+373.15 = 410.15K

z- Valancy of ion - for a monovalant cation z = +1

F - Faraday gas constant = 96485 C/mol

By applying into the equation, we get,

( (8.314 * 410.15) /(1 * 96485) ) / ln (80/10)

= 0.0353 * ln(8)

= 0.0353 * 2.079

= 0.0734V or 73.4mV

Therefore, the nerst potential for the given values are 73.4mV.